OpenStudy (anonymous):

If the range of the function f(x) = 4 − 4x is {-20, -16, -8, 0, 4}, what is its domain?

OpenStudy (anonymous):

Domain represents (x)..range represents f(x) or the solution..thus if you plug in each value for f(x) and solve for (x) you will have your domain values...

OpenStudy (anonymous):

can you help me by exampling the first range?

OpenStudy (anonymous):

For example, your first range is -20 plug this in the equation to find your domain (x) so it will be like this: f(x) = 4 − 4x -20 = 4 − 4x now, simplify this and solve for x. Then apply this method to get the corresponding domain for each given range in your question.